(-tan)^2/(1-cot)^2 = tan^2
1
Expert's answer
2019-11-05T12:05:47-0500

Solution:


"((-tan a)\/(1-cot a))^2-\u3016tan a\u3017^2=0"


"(tan a\/(cot a-1))^2-\u3016tan a\u3017^2=0"


"\u3016tan a\u3017^2 (1\/(cot a-1)^2 -1)=0"


"〖tan a〗^2=0 \,or \, 1/(cot a-1)^2 =1"

If "\u3016tan a\u3017^2=0" , then tan a=0, so the angle will be a=πn, n "\in" Z.

If "1\/(cot a-1)^2 =1" , then cot a =2, so the angle will be a=arccot 2+πm, m "\in" Z

                                 or cot a = 0, but in this situation tan a→∞

Answer:  πn, n "\in" Z; arccot 2+πm, m "\in" Z.